Find the general solution.

13. r (t) = = (5 -3) =(6) 5
Answer #1

13. 21- ZAx a = (>)< A=( A-1 - 1 5 -3 JA-111=0 =) 12+28 +2=0 3) =3512-2414-9--1£i 2 For tatti ! Av = (-1 tilv => V, -V2=(-1